OpenStudy (anonymous):

so confused 2/3n=4n-10/3n-1/2

OpenStudy (anonymous):

multiply both sides by 2

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

o sry is that 1/2

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

k then multiply both sides by 3n-1/2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1,5n

OpenStudy (anonymous):

=4n-10

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

1.5n=4n-10

OpenStudy (anonymous):

2.5/10 is it n=-4

OpenStudy (anonymous):

n=4

OpenStudy (anonymous):

because -10/-2.5

OpenStudy (anonymous):

2-1/n=4n-10 12=4n+n^-1

OpenStudy (anonymous):

4n^2+1=12n 4n^2-12n-1=0 quad